OpenStudy (anonymous):

A Ferrari accelerates from 0 to 100.0 km/h in 4 .80 s. What force (in Newtons) does a passenger of mass 96.0 kg experience during acceleration? I found; F=ma =556N ?

OpenStudy (anonymous):

100x(100/360)=27.7 (to change km/h to m/s) 27.7/4.8=5.787 (to solve acceleration) 5.787x96=555.5 (to solve force) therefore your answer is correct.

OpenStudy (anonymous):

We know that: \[F=ma\] We also know that: \[m=\frac{\Delta v}{\Delta t}=\frac{v_2-v_1}{t_2-t_1}\] So we can simplify: \[F=m\left(\frac{v_2-v_1}{t_2-t_1}\right)\] So then we have the following: \[\eqalign{ &m=96.0kg \\ &v_1=0m/s \\ &v_2=100.0km/h=\frac{100.0km}{h}\times\frac{1h}{3600s}\times\frac{1000m}{1km}=(100/36)m/s\approx27.78m/s \\ &t_1=0s \\ &t_2=4.8s \\ }\] So we can plug everything in: \[F=m\left(\frac{v_2-v_1}{t_2-t_1}\right)=(96.0)\left(\frac{27.78-0}{4.8-0}\right)=96\times\frac{27.78}{4.8}\approx555.6N\]

OpenStudy (anonymous):

Using significant digits though, you would round to 556N

OpenStudy (anonymous):

thanks.)

OpenStudy (anonymous):

Lol Anytime!